Zed is playing with a ball on top of a building but the ball fell and hits the ground after 2.6 seconds, what is the final velocity of the ball just before it hits the ground and how high is the building
1
Expert's answer
2018-03-12T16:36:07-0400
The displacement of the ball
h=(gt^2)/2
h=(9.8×〖2.6〗^2)/2=33.1 m
The final velocity of the ball
v=gt
v=9.8×2.6=25.5 m/s